# https://uni-fl.github.io/

### v1

@@ 4.1 Perceptual Feedback Learning
1:Dataset: Perceptual text-image dataset $\mathcal{D}=\{(\textrm{txt}_{1},\textrm{img}_{1}),...(\textrm{txt}_{n},\textrm{img}_{n})\}$
@@
2:Input: LDM with pre-trained parameters $w_{0}$, visual perceptual model $m_{\cdot}$, visual perceptual loss function $\Phi$, visual perceptual loss weight $\lambda$
@@
5:    $x_{0}$leftarrow VaeEnc($\textrm{img}_{i}$) // Obtain the latent representation of ground truth image
6:    $x_{T_{a}}$leftarrow AddNoise($x_{0}$) // Add noise into the latent according to Eq.1
@@
11:    $x^{{}^{\prime}}_{0}\leftarrow x_{t-1}$ // Predict the original latent by noise scheduler
@@
13:    $\mathcal{L}_{pefl}\leftarrow\lambda\Phi(m(\textrm{img}^{{}^{\prime}}_{i}),GT(\textrm{img}_{i})$ // PeFL loss by perceptual model
@@
Figure 3: The illustration of the PeFL on the (a) style and (b) structure optimization. The original noise MSE loss in the latent space only cares about the coarse reconstruction and overlooks the particular visual aspect of the generated image, which can be boosted by the existing various perceptual models via feedback learning.
4.2 Decoupled Feedback Learning
Decoupled Aesthetic Fine-tuning. Unlike objective visual quality, aesthetic quality is abstract and subjective, necessitating human aesthetic preference feedback to guide the model toward optimization based on human preferences. ImageReward  [57] addresses this issue via a human preference reward model trained on collected preference data within the ReFL framework. While effective, we argue that ImageReward is suboptimal as it relies on a single reward model trained with coarse annotated aesthetic preference data. The primary challenge arises from the attempt to encapsulate human preferences across multiple dimensions within a single reward model, which will result in inherent conflicts, as evidenced in certain Large Language Model (LLM) studies  [48]. To address this problem, we propose decoupling the different aesthetic dimensions during preference modeling to enable more effective aesthetic feedback learning. Specifically, we decompose the general aesthetic concept into representative dimensions and annotate them separately. These dimensions include color, layout, lighting, and detail. The data collection process is described in detail in the Appendix. Subsequently, we train an aesthetic preference reward model using this annotated data according to Eq.2. The objective of the decoupled feedback learning is:
$\mathcal{L}(\theta)_{aes}=\sum_{d}\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x^{\prime}_{0}\sim p(x^{\prime}_{0}|c)}[\texttt{ReLU}(\alpha_{d}-r_{d}(x^{\prime}_{0},c))]$ (6)
$r_{d}$ is the aesthetic reward model on $d$ dimension, $d\in\{\text{color},\text{layout},\text{detail},\text{lighting}\}$ and $\alpha_{d}$ are the dimension-aware hinge cofficient for loss calculation.
Active Prompt Selection. We observed that when using randomly selected prompts for preference fine-tuning, the diffusion model tends to rapidly overfit due to the limited semantic richness, leading to diminished effectiveness of the reward model. This phenomenon is commonly referred to as overoptimization  [62]. To address this issue, we further propose an active prompt selection strategy, which selects the most informative and diverse prompt from a prompt database. This selection process involves two key components: a semantic-based prompt filter and nearest neighbor prompt compression. By leveraging these techniques, the over-optimization can be greatly mitigated, achieving more efficient aesthetic reward fine-tuning. More details of this strategy are presented in the Appendix.
@@ 4.3 Adversarial Feedback Learning
The slow iterative denoising process employed in text-to-image diffusion models poses a significant hindrance to their practical application. To address this limitation, recent advancements, such as UFOGen  [58] and SDXL-Turbo  [42], have proposed incorporating adversarial training objectives into fine-tuning diffusion models. Building upon this insight, we introduce an adversarial feedback-learning method that combines feedback learning with the adversarial objective, aiming to accelerate the inference process.
The original optimization objective of the diffusion model seeks to increase the reward score of the output image, with the reward model held constant. Rather than freeze the reward model, we incorporate the optimization of an adversarial reward model $r_{a}(\cdot)$ during the fine-tuning process, treating it as a discriminator. In this way, the diffusion model serves as the generator and is optimized to enhance the reward score, while the reward model acts as the discriminator, aiming to distinguish between preferred and unpreferred samples. Consequently, the objective of adversarial feedback learning can be reformulated as follows:
@@
$\mathcal{L}^{D}(\phi)=-\mathbb{E}_{(x_{0},x^{\prime}_{0},c)\sim\mathcal{D}_{train},t\sim[1,T]}[\log\sigma(r_{a}(x_{0}))+\log(1-\sigma(r_{a}(x^{\prime}_{0})))].$
where $\theta$ and $\phi$ are the parameters of the diffusion model and discriminator. In practice, we follow PeFL to achieve adversarial training, considering the GT image as the preferred sample and the denoised image as the unpreferred sample. In this way, we continually guide the diffusion model to generate samples with higher fidelity and visual quality, which greatly accelerates the inference speed.
4.4 Training Pipeline
Our training process consists of two stages, each targeting a specific objective. In the first stage, we focus on improving visual generation quality and aesthetics. In the second stage, we apply adversarial feedback learning to accelerate the diffusion inference speed, which simultaneously updates the diffusion model and reward model with the adversarial training objective. We also integrate decoupled feedback learning to maintain aesthetics.
$\mathcal{L}^{1}(\theta)=\mathcal{L}(\theta)_{pefl}+\mathcal{L}(\theta)_{aes};\quad\mathcal{L}^{2}(\theta,\phi)=\mathcal{L}^{G}(\theta)+\mathcal{L}^{D}(\phi))+\mathcal{L}(\theta)_{aes}$ (8)
@@ 5.1 Implementation Details and Metrics
Dataset. For the PeFL training stage, we curated a large and high-quality dataset consisting of approximately 150k artist-style text images for style optimization and utilized COCO2017 [26] train split dataset with instance annotations and captions for structure optimization. Additionally, we collected the human preference dataset for the decoupled aesthetic feedback learning from diverse aspects (such as color, layout, detail, and lighting). The 100,000 prompts are selected for aesthetic optimization from DiffusionDB [52] through active prompt selection. During the adversarial feedback learning, we simply utilize an aesthetic subset of LAION [43] with image aesthetic scores above 5.
Training Setting. We employ the VGG-16 [45] network to extract gram matrix concepts for style PeFL and utilize the SOLO [51] as the instance segmentation model. We utilize the DDIM scheduler with a total of 20 inference steps. $T_{a}=10$ and the optimization steps $t\in[5,0]$ during PeFL training.
For adversarial feedback learning, we initialize the adversarial reward model with the weight of the aesthetic preference reward model of details. During adversarial training, the optimization step is set to $t\in[0,20]$ encompassing the entire process.
Baseline Models. We choose two representative text-to-image diffusion models with distinct generation capacities to comprehensively evaluate the effectiveness of UniFL, including (i) SD1.5 [36]; (ii) SDXL [32]. Based on these models, we pick up several state-of-the-art methods(i.e. ImageReward [57], Dreamshaper [9], and DPO [50] for generation quality improvement, LCM [27], SDXL-Turbo [40], and SDXL-Lightning [25] for inference acceleration) to compare the effectiveness of quality improvement and acceleration. All results of these methods are reimplemented with the official code provided by the authors.
Evaluation Metrics We generate the 5K image with the prompt from the COCO2017 validation split to report the Fréchet Inception Distance(FID) [20] as the overall visual quality metric. We also report the CLIP score with ViT-B-32 [12] and the aesthetic score with LAION aesthetic predictor to evaluate the text-to-image alignment and aesthetic quality of the generated images, respectively. Given the subjective nature of quality evaluations, we also conducted comprehensive user studies to obtain a more accurate evaluation. For more implementation details of UniFL, please refer to the Appendix.
Model Inference Steps FIDnormal-$\downarrow$ CLIP Scorenormal-$\uparrow$ Aestheticnormal-$\uparrow$
SD1.5 20 37.99 0.308 5.26
SD1.5 ImageReward [57] 20 32.31 0.312 5.37
SD1.5 DreamShaper [9] 20 34.21 0.313 5.44
SD1.5 DPO [50] 20 32.83 0.308 5.22
SD1.5 UniFL 20 31.14 0.318 5.54
SD1.5 4 42.91 0.279 5.16
SD1.5 LCM [27] 4 42.65 0.314 5.71
SD1.5 DreamShaper LCM [28] 4 35.48 0.314 5.58
SD1.5 UniFL 4 33.54 0.316 5.88
SDXL 25 27.92 0.321 5.65
SDXL ImageReward [57] 25 26.71 0.319 5.81
SDXL DreamShaper [9] 25 28.53 0.321 5.65
SDXL DPO [50] 25 35.30 0.325 5.64
SDXL UniFL 25 25.54 0.328 5.98
SDXL 4 125.89 0.256 5.18
SDXL LCM [27] 4 27.23 0.322 5.48
SDXL Turbo [40] 4 30.43 0.325 5.60
SDXL Lighting [25] 4 28.48 0.323 5.66
SDXL UniFL 4 26.25 0.325 5.87
Table 1: The quantitative comparison between our method and other methods on SD1.5 and SDXL architecture. The best performance is highlighted with bold font, and the second-highest performance is shown with underline.
@@ 5.2 Main Results
Quantitative Comparison. Tab.1 summarize the quantitative comparisons with competitive approaches on SD1.5 and SDXL. Generally, UniFL exhibits consistent performance improvement on both architectures and surpasses the existing methods of focus on improving generation quality or acceleration. Specifically, DreamShaper achieves considerable aesthetic quality in SD1.5(5.44), while ImageReard obtains the best performance in SDXL(5.88). Even though, UniFL surpasses these methods on all of these metrics in both SD1.5 and SDXL. In terms of acceleration, UniFL still exhibits notable performance advantages, surpassing the LCM with the same 4-step inference on both SD1.5 and SDXL. Surprisingly, we found that UniFL sometimes obtained even better aesthetic quality with fewer inference steps. For example, when applied to SD1.5, the aesthetic score is boosted from 5.26 to 5.54 without acceleration. After executing the acceleration with the adversarial feedback learning, the aesthetic score is further improved to 5.88 with much fewer inference steps. The related reasons will be investigated in the ablation experiment. We also compared the two latest acceleration methods on SDXL, including the SDXL Turbo and SDXL Lightning. Although retaining the high text-to-image alignment, we found that the image generated by SDXL Turbo tends to lack fidelity, leading to an inferior FID score. SDXL Lightning achieves the most balanced performance in all of these aspects and reaches impressive aesthetic quality in 4-step inference. However, UniFL still outperforms it in all kinds of metrics and achieves the best performance.
@@
User Study. We conducted a comprehensive user study using SDXL to evaluate the effectiveness of our method in enhancing generation quality and acceleration. As illustrated in Fig.4, our method significantly improves the original SDXL in terms of generation quality with 68% preference rate and outperforms DreamShaper and DPO by 36% and 25% preference rate, respectively. Thanks to perceptual feedback learning and decoupled aesthetic feedback learning, our method exhibits improvement even when compared to the competitive ImageReward, and is preferred by 17% additional people. In terms of acceleration, our method surpasses the widely used LCM by a substantial margin of 57% with 4-step inference. Even when compared to the latest acceleration methods like SDXL-Turbo and SDXL-Lightning, UniFL still demonstrates superiority and obtains more preference. This highlights the effectiveness of adversarial feedback learning in achieving acceleration.
@@
Figure 4: The user study about UniFL and other methods with 10 users on the generation of 500 prompts in generation quality(right) and inference acceleration(right).
@@
Figure 5: The visualization of the generation results of different methods based on SDXL.
@@
Qualitative Comparison As shown in Fig.5, UniFL achieves superior generation results compared with other methods. For example, when compared to ImageReward, UniFL generates images that exhibit a more coherent object structure (e.g., the horse), a more appropriate style (e.g., the cocktail), and a more captivating aesthetic quality (e.g., the warrior). Notably, even with fewer inference steps, UniFL consistently showcases higher generation quality, outperforming other methods. It is worth noting that SDXL-Turbo, due to its modification of the diffusion hypothesis, tends to produce images with a distinct style.

### v3

@@ 4.2 Decoupled Feedback Learning
1:  Dataset: Captioned perceptual text-image dataset with $\mathcal{D}=\{(\textrm{txt}_{1},\textrm{img}_{1}),...(\textrm{txt}_{n},\textrm{img}_{n})\}$
@@
2:  Input: LDM with pre-trained parameters $w_{0}$, perceptual model $m_{\cdot}$, perceptual loss function $\Phi$, loss weight $\lambda$
@@
5:     $x_{0}$leftarrow VaeEnc($\textrm{img}_{i}$) // From image to latent
6:     $x_{T_{a}}$leftarrow AddNoise($x_{0}$) // Add noise to latent
@@
11:     $x^{{}^{\prime}}_{0}\leftarrow x_{t-1}$ // Predict the denoised latent
@@
13:     $\mathcal{L}_{\text{pefl}}\leftarrow\lambda\Phi(m(\textrm{img}^{{}^{\prime}}_{i}),\mathrm{GT}(\textrm{img}_{i})$ // PeFL loss by perceptual model
@@
Active Prompt Selection. We observed that when using randomly selected prompts for aesthetic preference fine-tuning, the diffusion model tends to rapidly overfit the reward model due to the limited semantic richness, leading to diminished effectiveness of the reward model. To address this issue, we further propose an active prompt selection strategy, which selects the most informative and diverse prompt from a prompt database. This selection process involves two key components: a semantic-based prompt filter and nearest neighbor prompt compression. By leveraging these techniques, the overfitting can be greatly mitigated, achieving more efficient aesthetic reward fine-tuning. More details of this strategy are presented in the Appendix.B.2.
@@ 4.3 Adversarial Feedback Learning
The inherent iterative denoising process of diffusion models significantly hinders their inference speed. To address this limitation, we introduce adversarial feedback learning to reduce the denoising steps during inference. Specifically, to achieve inference acceleration, we exploit a general reward model $r_{a}(\cdot)$ to improve the generation quality of fewer denoising steps. However, as studied in [23], the samples under low inference steps tend to be too noisy to obtain the correct rewarding scores. To tackle this problem, rather than freeze the reward model during fine-tuning, we incorporate an extra adversarial optimization objective by treating $r_{a}(\cdot)$ as a discriminator and update it together with the diffusion model. Concretely, we follow a similar way with PeFL to take an image as input and execute the diffusion and denoising consecutively. Afterward, in addition to maximizing the reward score of the denoised image, we also update the reward model in an adversarial manner. The optimization objective is formulated as:
@@
$\mathcal{L}^{D}(\phi)=-\mathbb{E}_{(x_{0},x^{\prime}_{0},c)\sim\mathcal{D}_{\mathrm{train}},t\sim[1,T]}[\log\sigma(r_{a}(x_{0}))+\log(1-\sigma(r_{a}(x^{\prime}_{0})))].$
where $\theta$ and $\phi$ are the parameters of the diffusion model and discriminator. With the adversarial objective, the reward model is always aligned with the distribution of the denoised images with various denoised steps, enabling the reward model to function well across all the timesteps. Note that our method is distinct from the existing adversarial diffusion methods like SDXL-Turbo [24]. These methods take the adversarial distillation manner to accelerate the inference, which tends to require another LDM as the teacher model to realize distillation, incurring considerable memory costs. By contrast, we follow the reward feedback learning formulation, which integrates adversarial training with the reward tuning and achieves the adversarial reward feedback tuning via the lightweight reward model.
4.4 Training Objective
We employ a two-stage training pipeline to implement UniFL. The first stage focuses on improving generation quality, leveraging perceptual feedback learning and decoupled feedback learning to boost visual fidelity and aesthetic appeal. In the second stage, we apply adversarial feedback learning to accelerate the diffusion inference speed. To prevent potential degradation, we also include decoupled feedback learning to maintain aesthetics. The training objectives of each stage are summarized as follows:
$\mathcal{L}^{1}(\theta)=\mathcal{L}_{\mathrm{pefl}}(\theta)+\mathcal{L}_{\mathrm{aes}}(\theta);\quad\mathcal{L}^{2}(\theta,\phi)=\mathcal{L}^{G}(\theta)+\mathcal{L}^{D}(\phi)+\mathcal{L}_{\mathrm{aes}}(\theta)$ (7)
@@ 5.1 Implementation Details and Metrics
Dataset. We utilized the COCO2017 [47] train split dataset with instance annotations and captions for structure optimization with PeFL. Additionally, we collected the human preference dataset for the decoupled aesthetic feedback learning from diverse aspects (such as color, layout, detail, and lighting). 100,000 prompts are selected for aesthetic optimization from DiffusionDB [48] via active prompt selection. During the adversarial feedback learning, we use data from the aesthetic subset of LAION [49] with image aesthetic scores above 5.
Training Setting. We utilize the SOLO [50] as the instance segmentation model. We utilize the DDIM [51] scheduler with a total of 20 inference steps. $T_{a}=10$ and the optimization steps $t\in[0,5]$ during PeFL training. For adversarial feedback learning, we initialize the adversarial reward model with the weight of the aesthetic preference reward model of details. During adversarial training, the optimization step is set to $t\in[0,20]$ encompassing the entire diffusion process. Our training per stage costs around 200 A100 GPU hours.
Baseline Models. We choose two representative text-to-image diffusion models with distinct generation capacities to comprehensively evaluate the effectiveness of UniFL, including (i) SD1.5 [4]; (ii) SDXL [40]. Based on these models, we pick up several state-of-the-art methods(i.e. ImageReward [23], Dreamshaper [52], and DPO [22] for generation quality enhancement, LCM [27], SDXL-Turbo [24], and SDXL-Lightning [53] for inference acceleration) to compare the effectiveness of quality improvement and acceleration. All results of these methods are reimplemented with the official code provided by the authors.
Evaluation Metrics. We generate the 5K image with the prompt from the COCO2017 validation split to report the Fréchet Inception Distance (FID) [54] as the overall visual quality metric. We also report the CLIP score with ViT-B-32 [55] and the aesthetic score with LAION aesthetic predictor to evaluate the text-to-image alignment and aesthetic quality of the generated images, respectively. Given the subjective nature of quality evaluations, we further conducted comprehensive user studies to obtain a more accurate evaluation.
@@ 5.2 Main Results
Quantitative Comparison. Tab.1 summarize the quantitative comparisons with competitive approaches on SD1.5 and SDXL. Generally, UniFL exhibits consistent performance improvement on both architectures and surpasses the existing methods of focus on improving generation quality or acceleration. Specifically, for the generation quality, UniFL surpasses both DreamShaper (DS) and ImageReward (IR) across all metrics, where the former relies on high-quality training images while the latter exploits the human preference for fine-tuning. It is also the case when compared with the recently proposed preference tuning method DPO. In terms of acceleration, UniFL also exhibits notable performance advantages, surpassing the LCM with the same 4-step inference on both SD1.5 and SDXL. Surprisingly, we found that UniFL sometimes obtained even better aesthetic quality with fewer inference steps. For example, when applied to SD1.5, the aesthetic score is first boosted from 5.26 to 5.54 without acceleration, and then further improved to 5.88 after being optimized by adversarial feedback learning. This demonstrates the superiority of our method in acceleration.
[…]
SD1.5-Base 20 37.99 0.308 5.26
SD1.5-IR  [23] 20 32.31 0.312 5.37
SD1.5-DS [52] 20 34.21 0.313 5.44
SD1.5-DPO [22] 20 32.83 0.308 5.22
SD1.5-UniFL 20 31.14 0.318 5.54
SD1.5-Base 4 42.91 0.279 5.16
SD1.5-LCM [27] 4 42.65 0.314 5.71
SD1.5-DS LCM [26] 4 35.48 0.314 5.58
SD1.5-UniFL 4 33.54 0.316 5.88
SDXL-Base 25 27.92 0.321 5.65
SDXL-IR [23] 25 26.71 0.319 5.81
SDXL-DS [52] 25 28.53 0.321 5.65
SDXL-DPO [22] 25 35.30 0.325 5.64
SDXL-UniFL 25 25.54 0.328 5.98
SDXL-Base 4 125.89 0.256 5.18
SDXL-LCM [27] 4 27.23 0.322 5.48
SDXL-Turbo [24] 4 30.43 0.325 5.60
SDXL-Lighting [53] 4 28.48 0.323 5.66
SDXL-UniFL 4 26.25 0.325 5.87
Table 1: Quantitative comparison between our method and other methods on SD1.5 and SDXL architecture. The best performance is highlighted with bold font, and the second-best is underlined.
We also compared the two latest acceleration methods on SDXL, including the SDXL Turbo and SDXL Lightning. Although retaining the high text-to-image alignment, we found that the image generated by SDXL Turbo tends to lack fidelity, leading to an inferior FID score. SDXL Lightning achieves the most balanced performance in all of these aspects and reaches impressive aesthetic quality in 4-step inference. However, UniFL still obtains slightly better performance on these metrics.
@@
User Study. We conducted a comprehensive user study using SDXL to evaluate the effectiveness of our method in enhancing generation quality and acceleration. As illustrated in Fig.3, our method significantly improves the original SDXL in terms of generation quality with a 68% preference rate and outperforms DreamShaper and DPO by 36% and 25% preference rate, respectively. Thanks to PeFL and decoupled aesthetic feedback learning, our method exhibits improvement even when compared to the competitive ImageReward, and is preferred by 17% additional people. In terms of acceleration, our method surpasses the widely used LCM by a substantial margin of 57% with 4-step inference. Even when compared to the latest acceleration methods like SDXL-Turbo and SDXL-Lightning, UniFL still demonstrates superiority and obtains more preference. This highlights the effectiveness of adversarial feedback learning in achieving acceleration.
@@
Figure 3: User study about UniFL and other methods with 10 users on the generation of 500 prompts in generation quality (left) and inference acceleration (right).
@@
Figure 4: Qualitive comparison of the generation results of different methods based on SDXL.
@@
Qualitative Comparison. As shown in Fig.4, UniFL achieves superior generation results compared with other methods. For example, when compared to ImageReward, UniFL generates images that exhibit a more coherent object structure (e.g., the horse), and a more captivating aesthetic quality (e.g., the cocktail). Notably, even with fewer inference steps, UniFL consistently showcases higher generation quality, outperforming other methods. It is worth noting that SDXL-Turbo, due to its modification of the diffusion hypothesis, tends to produce images with a distinct style.